OpenStudy (anonymous):

Convert the mixed number, 1 3/4 , to a fraction. Reduce fraction if necessary. A. 4/12 B. 7/4 C. 4/7 D. 12/4

terenzreignz (terenzreignz):

Okay, there are three parts of your mixed number, right? The whole number, the numerator, and the denominator.

OpenStudy (stamp):

see attachment

OpenStudy (anonymous):

not bad @stamp

terenzreignz (terenzreignz):

WHOLE NUMBER NUMERATOR DENOMINATOR To convert from mixed-number to a fraction, get the denominator, multiply it to the whole number, then add that product to the numerator. THAT is your numerator, and just maintain the denominator. To illustrate \[\huge a\frac{b}{c}=\frac{c\times a+b}{c}\]

OpenStudy (anonymous):

so it would be 7/4?

OpenStudy (stamp):

...

OpenStudy (anonymous):

look at how stamp worked it out in the attachment

terenzreignz (terenzreignz):

LOL @stamp already showed it explicitly :) But let's do it... step-by-step WHOLE NUMBER is 1 NUMERATOR is 3 DENOMINATOR is 4 So... four times 1 is 4, plus 3, is 7, that's your numerator your denominator remains... 4

OpenStudy (anonymous):

ok. so the bottom number will be 4

terenzreignz (terenzreignz):

Yes. Don't forget to reduce to lowest terms, if possible, but 7/4 is already reduced.

OpenStudy (anonymous):

ok thanks